Word problem: Find two consecutive even integers such that the smaller integer subtracted from twice the larger gives a difference of 10. find the two even integers.
I go the answers just by trial and error as: 6 & 8 but is there a formula to use for this problem? thanks

Found 3 solutions by Mathtut, Earlsdon, ankor@dixie-net.com:
Answer by Mathtut(3670)

2 CONSECUTIVE EVEN integers. that woud make the number 2 apart from each other.
:
lets call one integer,the smaller, a and the other b. But we know that b is equal to a+2
:
2(a+2)-a=10
:
2a+4-a=10
:
a=6
:
b=a+2=8
:
there ya go

Answer by Earlsdon(6294)

Let x = the smaller even integer and x+2 = the next consecutive even integer.
2(x+2)-x = 10 "...the smaller integer subtracted from twice the larger integer gives a difference of 10."
2x+4-x = 10
x+4 = 10
x = 6 and x+2 = 8
The two consecutive even integers are 6 and 8.

Answer by ankor@dixie-net.com(22740)

Find two consecutive even integers such that the smaller integer subtracted from twice the larger gives a difference of 10. find the two even integers.
:
Here's an easy way, write an equation for the statement
:
let the two consecutive integers be: x, (x+2)
:
"smaller integer subtracted from twice the larger gives a difference of 10.
2(x + 2) - x = 10
2x + 4 - x = 10
2x - x = 10 - 4
x = 6 and 8 are the integers